# On graphs with large third eigenvalue

### v3

@@ 1 Introduction
This gives that $c_{2}=\frac{1}{2}$, as the construction $kK_{\frac{n}{k}}$ has $\lambda_{k}(kK_{\frac{n}{k}})=\frac{n}{k}-1$, so $c_{k}\geq\frac{1}{k}$. Yet for $k\geq 5$, Nikiforov provided a construction yielding $c_{k}>\frac{1}{k-\frac{1}{2}}$. Additionally, for sufficiently large $k$, Nikiforov used the family of strongly-regular graphs introduced by Taylor in [11, 12], widely known as Taylor graphs, for a closely matching lower bound to Theorem 1.1.
@@
There exists $k_{0}$ such that if $k>k_{0}$, then
@@
We then prove that multiple families of graphs, which provide the best known lower bounds for larger values of $k$, have $\lambda_{3}(G)\leq\frac{|V(G)|}{3}$, further supporting Question 1.4. In particular, we prove and generalise the following theorem.
@@
If there exists a minimiser of $\lambda_{n-1}$ over $\mathcal{S}_{n}$ with simple final eigenvalue, then there exists an $11\times 1$ 11 matrix $T$ with entries in $\{0,1\}$ and a vertex multiplication $\tilde{T}$ of $T$ with:
@@
A closed vertex multiplication of a graph $G$ is constructed by replacing each vertex of $G$ with a clique and each edge of $G$ with a complete bipartite graph between the endpoint cliques. If the cliques are the same size, this reduces to the closed blow-up.
@@ 3.1 Strongly regular graphs
We make use of the Theorem 3.6.7 from Cvetković, Rowlinson and Simić [5], stated as a lemma here:
@@
If $G$ is disconnected, then $f=0$. As $r(r-e-1)=(n-r-1)f$ for strongly regular graphs, either $r=0$ or $r=e+1$. In the former case, $G\cong E_{n}$ has $\lambda_{3}(G)=0<\frac{n}{3}$. In the latter case, we have $r+1\mid n$ and $G\cong\frac{n}{r+1}K_{r+1}$, so $G$ has spectrum $r^{n/(r+1)},(-1)^{nr/(r+1)}$. For $\lambda_{3}(G)\geq\frac{n}{3}$, we require $\frac{n}{r+1}\geq 3$, but this gives $\lambda_{3}(G)=r\leq\frac{n}{3}-1<\frac{n}{3}$. Similarly, if $\bar{G}$ is disconnected, then $G$ must be the $m$-fold blow-up of the complete graph $K_{n/m}$, where $m\mid n$. Thus, $G$ has spectrum $m(\frac{n}{m}-1),0^{\left(n-\frac{n}{m}\right)},(-m)^{\left(\frac{n}{m}-1\right)}$. So, $\lambda_{3}(G)=0<\frac{n}{3}$.
@@ 3.2 Regular line graphs
Then, note that $BB^{T}=M+rI$ and $B^{T}B=N+2I$. As $BB^{T}$ and $B^{T}B$ share non-zero eigenvalues, we have $\chi_{N}(t)=(t+2)^{m-n}\chi_{M}(t-r+2)$. Thus, $\lambda_{3}(N)$ is either $-2$ or $\lambda_{3}(M)+r-2$. In any case
@@
If $n_{2}=4$, we are done unless $r_{2}<3r_{1}-6$. As $r_{1}\leq 4$, we must have $r_{2}\leq 5$. Moreover, $r_{2}\leq 3r_{2}-6$ implies that $r_{2}\geq 3$. There are only 7 possible sets of parameters: $(5,4,4,5),(4,4,4,4),(8,4,2,4),(16,4,1,4),(4,4,3,3),(6,4,2,3),(12,4,1,3)$ and there are only 58 graphs with these parameters. We may check these manually and yield that each such line graph has third eigenvalue proportion $\leq 0.202254248$.
If $n_{2}=3$, then either $G\cong 3K_{1,r_{2}}$, $G\cong K_{3,n_{1}}$ or $G$ is isomorphic to the closed vertex multiplication of $C_{6}$ by $[1,r_{2},1,r_{2},1,r_{2}]$. In any case, we have $\frac{\lambda_{3}(L(G))}{|V(L(G))|}\leq\frac{1}{3}$.
If $n_{2}=2$, we immediately have either $G\cong 2K_{1,r_{2}}$ or $G\cong K_{2,n_{1}}$. In the former case, we have $\lambda_{3}(L(G))=\lambda_{3}(2K_{r_{2}})=-1$ and in the latter, $\lambda_{3}(L(G))=\lambda_{3}(K_{n_{1}}\square P_{2})=0$. Finally, if $n_{2}=1$ we have $G\cong K_{1,n_{1}}$. ∎
@@ 3.3 Cayley graphs on abelian groups
Consider the $n$-th roots of unity on the complex plane. Then the ones with non-negative real part are those with argument in $[-\pi/2,\pi/2]$ hence correspond exactly to $\{\omega^{-\lfloor{n/4\rfloor}},\omega^{-\lfloor{n/4\rfloor}+1},\dots,\omega^{0},\dots,\omega^{\lfloor{n/4\rfloor}}\}$. Since $\omega^{k}$ and $\omega^{-k}$ are complex conjugates, $\omega^{k}+\omega^{-k}$ is real and the sum of this subset is real, which by construction attains maximal real part. ∎
@@
Consider $\delta_{n}=f(n)-\cot(\pi/n)=|\frac{\omega^{2\lfloor{n/4\rfloor}+1}-1}{\omega-1}|-|\frac{\omega^{n/2+1}-1}{\omega-1}|$. By the triangle inequality, $|\delta_{n}|\leq\frac{|\omega^{2\lfloor{n/4\rfloor}-n/2}-1|}{|\omega-1|}$. Note that $2\lfloor{n/4\rfloor}-n/2\in\{0,-1/2,-1,-3/2\}$. Hence for $n>150$, by considering the distances between roots of unity on one quarter-arc of a circle, we get $|\omega^{2\lfloor{n/4\rfloor}-n/2}-1|\leq|\omega^{2}-1|\leq|\omega^{2}-\omega|+|\omega-1|=2|\omega-1|$. Thus, $|\delta_{n}|\leq 2$, hence we have $\frac{f(n)}{n}\leq\frac{\cot(\pi/n)+2}{n}$ and
@@
The function $\frac{\cot(\pi/x)+2}{x}$ has derivative $\frac{\pi\cot(\pi/x)^{2}+\pi-x\cot(\pi/x)-2x}{x^{3}}$. Note further that $\tan(x)>x$ for $x\in(0,\pi/2)$, hence $\cot(\pi/x)<x/\pi$ for $x>2$. Now $\pi y^{2}-xy=y(\pi y-x)<0$ for $0<y<x/\pi$, hence in particular $\pi\cot(\pi/x)^{2}+\pi-x\cot(\pi/x)-2x<\pi-2x<0$ for $x>2$ and our function is decreasing. We compute for $x=150$ the value $0.3315...<1/3$, hence $f(n)<n/3$ for $n>150$ as well. ∎
@@
Eigenvalues of graphs are real. By partitioning the multiset $\chi(S)$ into $d$ subsets of distinct elements, each subset has real part at most $f(k)$, hence
@@
In particular, $n/4$ elements of $\Gamma$ have $(\chi_{1}(g),\chi_{2}(g))=(+1,+1)$, and identically for $(+1,-1)$, $(-1,+1)$, $(-1,-1)$. Thus, $\lambda_{\chi_{1}}+\lambda_{\chi_{2}}=\sum_{s\in S}\chi_{1}(s)+\chi_{2}(s)\leq n/2$ and by pigeonhole principle, one of $\lambda_{\chi_{1}},\lambda_{\chi_{2}}\leq n/4$. In particular, at most one character with $k=2$ can have $\lambda>n/4$, so $\lambda_{3}(G)\leq n/3$ always. ∎
@@ 3.4 Semi-Cayley graphs on abelian groups
Let $\Gamma$ be a group and $R,S,T\subset\Gamma$ with $R=R^{-1},S=S^{-1}$ and $e\not\in R\cup S$. The semi-Cayley graph $SC(\Gamma;R,S,T)$ is the graph with $V(G)=\Gamma\times\{0,1\}$ with $(h,i)\sim(g,j)$ if and only if one of the following holds: {outline} \1 $i=j=0$ and $gh^{-1}\in R$, \1 $i=j=1$ and $gh^{-1}\in S$, \1 $i=0,j=1$ and $gh^{-1}\in T$. Using similar techniques to above, we prove the following result.
@@
For $k\geq 3$ and $4\nmid k$, this gives $\frac{\lambda_{\chi}}{2n}\leq\frac{f(k)}{k}\leq\frac{1}{3}$ as before. For $k\geq 3$ and $4\mid k$ note that we get the equality $02\lfloor{n/4\rfloor}-n/2=$ 0 hence $\delta_{n}=0$ in the proof of Lemma 3.8. Thus $f(k)=\cot(\pi/k)\leq k/\pi$. In particular, $\frac{f(k)+g(k)}{2k}\leq\frac{k/\pi+\sqrt{(k/\pi)^{2}+1}}{2k}=\frac{1+\sqrt{1+(\pi/k)^{2}}}{2\pi}$. This is a decreasing function in $k$ and is less than $1/3$ for $k=8$. By direct computation, this also holds for $k=4$, hence for $k\geq 3$ in general we again obtain $\frac{\lambda_{\chi}}{2n}\leq\frac{1}{3}$.
To control $k=2$, we first bound $\lambda_{\chi}$ in two cases: {outline}[enumerate] \1 Case 1: $|\lambda_{\chi}^{T}|\leq n/6$. Then note that $(\lambda_{\chi}^{R}-\lambda_{\chi}^{S})^{2}+4|\lambda_{\chi}^{T}|^{2}\leq(|\lambda_{\chi}^{R}-\lambda_{\chi}^{S}|+2|\lambda_{\chi}^{T}|)^{2}$ hence
@@
We also claim that $|\lambda_{\chi_{1}}^{T}|+|\lambda_{\chi_{2}}^{T}|\leq n/2$. Indeed, suppose there are $p$ elements of $T$ satisfying $(\chi_{1}(t),\chi_{2}(t))=(+1,+1)$. Analogously, define $q,r,s$ for $(+1,-1),(-1,+1),(-1,-1)$ respectively. Then we get $|p+q-r-s|+|p-q+r-s|=2u-2v\leq 2u\leq n/2$ for some $u,v\subset\{p,q,r,s\}$, where $40\leq p,q,r,s\leq n/$ 4. Thus, if we let $|\lambda_{\chi_{1}}^{T}|=y$, we have $|\lambda_{\chi_{2}}^{T}|\leq n/2-y$.
As before, we prove that at least one of $k=2$ gives $\lambda_{\chi}\leq\frac{2n}{3}$ via pigeonhole. Note that the smaller eigenvalue for each $\chi$ is bounded above by $\frac{n/2+n/2}{2}=n/2$, hence it suffices to consider only the larger one. {outline} \1[-] If $|\lambda_{\chi}^{R}-\lambda_{\chi}^{S}|\leq n/2$ holds for both, we obtain
@@
Regarding as a function of $y$, we get a maximum at $y=0$ (or $y=n/2$) where we get $\frac{3+\sqrt{5}}{4}n<\frac{4n}{3}$. \1[-] If $|\lambda_{\chi}^{R}-\lambda_{\chi}^{S}|\geq n/2$ holds for both, then we either have both satisfying $|\lambda_{\chi}^{T}|\leq n/3$, or one yields $|\lambda_{\chi}^{T}|>n/3$ and the other $|\lambda_{\chi}^{T}|<n/6$, since $|\lambda_{\chi_{1}}^{T}|+|\lambda_{\chi_{2}}^{T}|\leq n/2$. In the former case, both are Case 2, and in the latter, we have one of each of Case 1, Case 2. Thus $\lambda_{\chi_{1}}+\lambda_{\chi_{2}}\leq\frac{4n}{3}$. \1[-] If one of each, suppose WLOG $|\lambda_{\chi_{1}}^{R}-\lambda_{\chi_{1}}^{S}|=n/2+t>n/2$. Then as in Case 2 we have say $\lambda_{\chi_{1}}^{R}\geq 0$ and $\lambda_{\chi_{1}}^{S}\leq-t$, hence $\lambda_{\chi_{1}}^{S}+\lambda_{\chi_{2}}^{S}\leq n/2-t$. Combined with $\lambda_{\chi_{1}}^{R}+\lambda_{\chi_{2}}^{R}\leq n/2$, we get
@@
over $20\leq t\leq n/$ 2, $20\leq y\leq n/$ 2. The derivative with respect to $t$ is $\frac{1}{2}(-1+\frac{n/2+t}{\sqrt{(n/2+t)^{2}+4y^{2}}})$ which is always non-positive. Thus, maximum lies at $t=0$, reducing to (*) where $\lambda_{\chi_{1}}+\lambda_{\chi_{2}}\leq\frac{3+\sqrt{5}}{4}n<\frac{4n}{3}$.
The final case to eliminate is if both eigenvalues for $k=1$, which are $\frac{|R|+|S|\pm\sqrt{(|R|-|S|)^{2}+4|T|^{2}}}{2}$ and one for $k=2$ yield $\lambda_{\chi}>\frac{2n}{3}$. Let $\chi$ correspond to $k=2$. Let $\frac{|R|+|S|}{2}=\frac{2n}{3}+y$, then smaller eigenvalue greater than $\frac{2n}{3}$ means $\sqrt{(|R|-|S|)^{2}+4|T|^{2}}<2y$ and $|T|<y$. Note also $|R|+|S|\leq 2n$ so $y\leq n/3$. If we further have $y\leq n/6$ then this falls in Case 1, else if $n/6\leq y\leq n/3$ with $|\lambda_{\chi}^{R}-\lambda_{\chi}^{S}|\geq n/2$ we get Case 2. Otherwise, $y>n/6$ and $|\lambda_{\chi}^{R}-\lambda_{\chi}^{S}|<n/2$.
Since there are $n/2$ of each of ±$\pm 1$ in $\chi(\Gamma)$, if $|R|=n/2+r$ then necessarily $\lambda_{\chi}^{R}\leq n/2-r$ and similarly for $S$. In particular, $|R|+|S|=\frac{4n}{3}+2y$ and therefore $\lambda_{\chi}^{R}+\lambda_{\chi}^{S}\leq n-(n/3+2y)=\frac{2n}{3}-2y$. Hence altogether we have

### v4

@@ 1 Introduction
This gives that $c_{2}=\frac{1}{2}$, as the construction $kK_{\frac{n}{k}}$ has $\lambda_{k}(kK_{\frac{n}{k}})=\frac{n}{k}-1$, so $c_{k}\geq\frac{1}{k}$. Nikiforov provided a further construction yielding $c_{k}\geq\frac{1}{k-\frac{1}{2}}$ for $k\geq 5$. Additionally, for sufficiently large $k$, Nikiforov used the family of strongly-regular graphs introduced by Taylor in [11, 12], widely known as Taylor graphs, for a closely matching lower bound to Theorem 1.1.
@@
There exists a $k_{0}$ such that if $k>k_{0}$, then
@@
We then prove that multiple families of graphs, which provide the best known lower bounds for larger values of $k$, have $\lambda_{3}(G)\leq\frac{|V(G)|}{3}$, further suggesting $c_{3}=\frac{1}{3}$. In particular, we prove and generalise the following theorem.
@@
If there exists a minimiser of $\lambda_{n-1}$ over $\mathcal{S}_{n}$ with simple final eigenvalue, then there exists an $11\times 11$ matrix $T$ with entries in $\{0,1\}$ and a vertex multiplication $\tilde{T}$ of $T$ with:
@@
A closed vertex multiplication of a graph $G$ is constructed by replacing each vertex of $G$ with a clique and each edge of $G$ with a complete bipartite graph between the endpoint cliques. For a graph $G$ with vertices $v_{1},\dots,v_{n}$, the closed vertex multiplication of $G$ by $a_{1},\dots,a_{n}]$ is the closed vertex multiplication where the vertex $v_{i}$ is replaced by a clique of size $a_{i}$. If the cliques are chosen to be the same size, this reduces to a closed blow-up.
@@ 3.1 Strongly regular graphs
We make use of the absolute bound for strongly regular graphs, which is given as Theorem 3.6.7 from Cvetković, Rowlinson and Simić [5] and stated as a lemma here:
@@
If $G$ is disconnected, then $f=0$. As r(re1)=(nr1)fr(r-e-1)=(n-r-1)f for strongly regular graphs, either $r=0$ or $r=e+1$. In the former case, $G\cong\overline{K_{n}}$ has $\lambda_{3}(G)=0<\frac{n}{3}$. In the latter case, we have $r+1\mid n$ and $G\cong\frac{n}{r+1}K_{r+1}$, so $G$ has spectrum $r^{n/(r+1)},(-1)^{nr/(r+1)}$. For $\lambda_{3}(G)\geq\frac{n}{3}$, we require $\frac{n}{r+1}\geq 3$, but this gives $\lambda_{3}(G)=r\leq\frac{n}{3}-1<\frac{n}{3}$. Similarly, if $\bar{G}$ is disconnected, then $G$ must be the $m$-fold blow-up of the complete graph $K_{n/m}$, where $m\mid n$. Thus, $G$ has spectrum $m(\frac{n}{m}-1),0^{\left(n-\frac{n}{m}\right)},(-m)^{\left(\frac{n}{m}-1\right)}$. So, $\lambda_{3}(G)=0<\frac{n}{3}$.
@@ 3.2 Regular line graphs
Then, note that $BB^{T}=M+rI$ and $B^{T}B=N+2I$. As $BB^{T}$ and $B^{T}B$ share non-zero eigenvalues, we have $\chi_{N}(t)=(t+2)^{m-n}\chi_{M}(t-r+2)$. Thus, $\lambda_{3}(N)$ is either 2-2 or $\lambda_{3}(M)+r-2$. In any case
@@
If $n_{2}=4$, we are done unless $r_{2}<3r_{1}-6$. As $r_{1}\leq 4$, we must have $r_{2}\leq 5$. Moreover, $r_{2}\leq 3r_{2}-6$ implies that $r_{2}\geq 3$. There are only 7 possible sets of parameters: (5,4,4,5),(4,4,4,4),(8,4,2,4),(16,4,1,4),(4,4,3,3),(6,4,2,3),(12,4,1,3)(5,4,4,5),(4,4,4,4),(8,4,2,4),(16,4,1,4),(4,4,3,3),(6,4,2,3),(12,4,1,3) and there are only 58 graphs with these parameters. We may check these manually and yield that each such line graph has third eigenvalue proportion $\leq 0.202254248$.
If $n_{2}=3$, then either $G\cong 3K_{1,r_{2}}$, $G\cong K_{3,n_{1}}$ or $G$ is isomorphic to the closed vertex multiplication of $C_{6}$ by $1,r_{2},1,r_{2},1,r_{2}]$. In any case, we have $\frac{\lambda_{3}(L(G))}{|V(L(G))|}\leq\frac{1}{3}$.
If $n_{2}=2$, we immediately have either $G\cong 2K_{1,r_{2}}$ or $G\cong K_{2,n_{1}}$. In the former case, we have $\lambda_{3}(L(G))=\lambda_{3}(2K_{r_{2}})=-1$ and in the latter, $\lambda_{3}(L(G))=\lambda_{3}(K_{n_{1}}\square P_{2})=0$. Finally, if $n_{2}=1$ we have $G\cong K_{1,n_{1}}$ and so $\lambda_{3}(L(G))=-1$. In any case, we have that $\lambda_{3}(L(G))=\lambda_{3}(K_{n_{1}})\leq\frac{|V(L(G))|}{3}$, completing the proof. ∎
@@ 3.3 Cayley graphs on abelian groups
Consider the $n$-th roots of unity on the complex plane. Then the ones with non-negative real part are those with argument in $\pi/2,\pi/2]$ hence correspond exactly to $\{\omega^{-\lfloor{n/4\rfloor}},\omega^{-\lfloor{n/4\rfloor}+1},\dots,\omega^{0},\dots,\omega^{\lfloor{n/4\rfloor}}\}$. Since $\omega^{k}$ and $\omega^{-k}$ are complex conjugates, $\omega^{k}+\omega^{-k}$ is real and the sum of this subset is real, which by construction attains maximal real part. ∎
@@
Consider $\delta_{n}=f(n)-\cot(\pi/n)=|\frac{\omega^{2\lfloor{n/4\rfloor}+1}-1}{\omega-1}|-|\frac{\omega^{n/2+1}-1}{\omega-1}|$. By the triangle inequality, $|\delta_{n}|\leq\frac{|\omega^{2\lfloor{n/4\rfloor}-n/2}-1|}{|\omega-1|}$. Note that $\alpha:=2\lfloor{n/4\rfloor}-n/2\in\{0,-1/2,-1,-3/2\}$. Since $n\geq 4$ and $|\sin(x)|$ is a decreasing function of $x$ in $\pi/2,0]$, we have:
$|\omega^{2\lfloor{n/4\rfloor}-n/2}-1|=2|\sin(\pi\alpha/n)|\leq 2|\sin(-2\pi/n)|=|\omega^{2}-1|.$
Applying the triangle inequality again, we get $|\omega^{2\lfloor{n/4\rfloor}-n/2}-1|\leq|\omega^{2}-\omega|+|\omega-1|=2|\omega-1|$. Thus, $|\delta_{n}|\leq 2$, hence we have $\frac{f(n)}{n}\leq\frac{\cot(\pi/n)+2}{n}$ and
@@
The function $\frac{\cot(\pi/x)+2}{x}$ has derivative $\frac{\pi\cot(\pi/x)^{2}+\pi-x\cot(\pi/x)-2x}{x^{3}}$. Note further that $\tan(x)>x$ for $x\in(0,\pi/2)$, hence $\cot(\pi/x)<x/\pi$ for $x>2$. Now $\pi y^{2}-xy=y(\pi y-x)<0$ for $0<y<x/\pi$, hence in particular $\pi\cot(\pi/x)^{2}+\pi-x\cot(\pi/x)-2x<\pi-2x<0$ for $x>2$ and our function is decreasing. We compute for $x=150$ the value 0.3315<1/30.3315...<1/3, hence $f(n)<n/3$ for $n>150$ as well. ∎
@@
Eigenvalues of graphs are real. We construct a partition of the multiset $\chi(S)$ into $d$ subsets of distinct elements as follows: for each element in the set $\chi(S)$, add its $i$-th copy to subset $i$. By definition of $d$, we end up with exactly $d$ subsets and each subset has real part at most $f(k)$, hence:
@@
In particular, $n/4$ elements of $\Gamma$ have $(\chi_{1}(g),\chi_{2}(g))=(+1,+1)$, and identically for (+1,1)(+1,-1), (1,+1)(-1,+1), (1,1)(-1,-1). Thus, $\lambda_{\chi_{1}}+\lambda_{\chi_{2}}=\sum_{s\in S}\chi_{1}(s)+\chi_{2}(s)\leq n/2$ and by pigeonhole principle, one of $\lambda_{\chi_{1}},\lambda_{\chi_{2}}\leq n/4$. In particular, at most one character with $k=2$ can have $\lambda>n/4$, so $\lambda_{3}(G)\leq n/3$ always. ∎
@@ 3.4 Semi-Cayley graphs on abelian groups
Let $\Gamma$ be a group and $R,S,T\subset\Gamma$ with $R=R^{-1},S=S^{-1}$ and $e\not\in R\cup S$. The semi-Cayley graph $SC(\Gamma;R,S,T)$ is the graph with $V(G)=\Gamma\times\{0,1\}$ with $(h,i)\sim(g,j)$ if and only if one of the following holds:
  • $i=j=0$ and $gh^{-1}\in R$,
  • $i=j=1$ and $gh^{-1}\in S$,
  • $i=0,j=1$ and $gh^{-1}\in T$.
Using similar techniques to above, we prove the following result.
@@
For $k\geq 3$ and $4\nmid k$, this gives $\frac{\lambda_{\chi}}{2n}\leq\frac{f(k)}{k}\leq\frac{1}{3}$ as before. For $k\geq 3$ and $4\mid k$ note that we get the equality $2\lfloor{n/4\rfloor}-n/2=0$ hence $\delta_{n}=0$ in the proof of Lemma 3.8. Thus $f(k)=\cot(\pi/k)\leq k/\pi$. In particular, $\frac{f(k)+g(k)}{2k}\leq\frac{k/\pi+\sqrt{(k/\pi)^{2}+1}}{2k}=\frac{1+\sqrt{1+(\pi/k)^{2}}}{2\pi}$. This is a decreasing function in $k$ and is less than $1/3$ for $k=8$. By direct computation, this also holds for $k=4$, hence for $k\geq 3$ in general we again obtain $\frac{\lambda_{\chi}}{2n}\leq\frac{1}{3}$.
To control $k=2$, we first bound $\lambda_{\chi}$ in two cases:
  1. 1.
    Case 1: $|\lambda_{\chi}^{T}|\leq n/6$. Then note that $(\lambda_{\chi}^{R}-\lambda_{\chi}^{S})^{2}+4|\lambda_{\chi}^{T}|^{2}\leq(|\lambda_{\chi}^{R}-\lambda_{\chi}^{S}|+2|\lambda_{\chi}^{T}|)^{2}$ hence
@@
We also claim that $|\lambda_{\chi_{1}}^{T}|+|\lambda_{\chi_{2}}^{T}|\leq n/2$. Indeed, suppose there are $p$ elements of $T$ satisfying $(\chi_{1}(t),\chi_{2}(t))=(+1,+1)$. Analogously, define $q,r,s$ for (+1,1),(1,+1),(1,1)(+1,-1),(-1,+1),(-1,-1) respectively. Then we get $|p+q-r-s|+|p-q+r-s|=2u-2v\leq 2u\leq n/2$ for some $u,v\subset\{p,q,r,s\}$, where $0\leq p,q,r,s\leq n/4$. Thus, if we let $|\lambda_{\chi_{1}}^{T}|=y$, we have $|\lambda_{\chi_{2}}^{T}|\leq n/2-y$.
As before, we prove that at least one of $\chi_{1},\chi_{2}$ gives $\lambda_{\chi}\leq\frac{2n}{3}$ via pigeonhole. Note that the smaller eigenvalue for each $\chi$ is bounded above by $\frac{n/2+n/2}{2}=n/2$, hence it suffices to consider only the larger one.
    If $|\lambda_{\chi}^{R}-\lambda_{\chi}^{S}|\leq n/2$ holds for both $\chi_{1},\chi_{2}$, we obtain
@@
    Regarding the right-hand side of (*) as a function of $y$, we get a maximum at $y=0$ (or $y=n/2$) where we get $\frac{3+\sqrt{5}}{4}n<\frac{4n}{3}$.
    If $|\lambda_{\chi}^{R}-\lambda_{\chi}^{S}|\geq n/2$ holds for both, then we either have both satisfying $|\lambda_{\chi}^{T}|\leq n/3$, or one yields $|\lambda_{\chi}^{T}|>n/3$ and the other $|\lambda_{\chi}^{T}|<n/6$, since $|\lambda_{\chi_{1}}^{T}|+|\lambda_{\chi_{2}}^{T}|\leq n/2$. In the former case, by Case 2 we have $\lambda_{\chi_{1}},\lambda_{\chi_{2}}\leq\frac{2n}{3}$, and in the latter case, by Case 1 and Case 2 we again have $\lambda_{\chi_{1}},\lambda_{\chi_{2}}\leq\frac{2n}{3}$. Thus $\lambda_{\chi_{1}}+\lambda_{\chi_{2}}\leq\frac{4n}{3}$.
    If one of each, suppose WLOG $|\lambda_{\chi_{1}}^{R}-\lambda_{\chi_{1}}^{S}|=n/2+t>n/2$. Then as in Case 2 we have say $\lambda_{\chi_{1}}^{R}\geq 0$ and $\lambda_{\chi_{1}}^{S}\leq-t$, hence $\lambda_{\chi_{1}}^{S}+\lambda_{\chi_{2}}^{S}\leq n/2-t$. Combined with $\lambda_{\chi_{1}}^{R}+\lambda_{\chi_{2}}^{R}\leq n/2$, we get
@@
    over $0\leq t\leq n/2$, $0\leq y\leq n/2$. The derivative with respect to $t$ is $\frac{1}{2}(-1+\frac{n/2+t}{\sqrt{(n/2+t)^{2}+4y^{2}}})$ which is always non-positive. Thus, the maximum lies at $t=0$, reducing to (*) where $\lambda_{\chi_{1}}+\lambda_{\chi_{2}}\leq\frac{3+\sqrt{5}}{4}n<\frac{4n}{3}$.
The final case to eliminate is if both eigenvalues for $k=1$, which are $\frac{|R|+|S|\pm\sqrt{(|R|-|S|)^{2}+4|T|^{2}}}{2}$ and one for $k=2$ yield $\lambda_{\chi}>\frac{2n}{3}$. Let $\chi$ correspond to $k=2$. Let $\frac{|R|+|S|}{2}=\frac{2n}{3}+y$, then since $\lambda_{\chi}>\frac{2n}{3}$, $\sqrt{(|R|-|S|)^{2}+4|T|^{2}}<2y$ and $|T|<y$. Note also $|R|+|S|\leq 2n$ so $y\leq n/3$. If we further have $y\leq n/6$ then this falls in Case 1, else if $n/6\leq y\leq n/3$ with $|\lambda_{\chi}^{R}-\lambda_{\chi}^{S}|\geq n/2$ we get Case 2. Otherwise, $y>n/6$ and $|\lambda_{\chi}^{R}-\lambda_{\chi}^{S}|<n/2$.
Since there are $n/2$ of each of $\pm 1$ in $\chi(\Gamma)$, if $|R|=n/2+r$ then necessarily $\lambda_{\chi}^{R}\leq n/2-r$ and similarly for $S$. In particular, $|R|+|S|=\frac{4n}{3}+2y$ and therefore $\lambda_{\chi}^{R}+\lambda_{\chi}^{S}\leq n-(n/3+2y)=\frac{2n}{3}-2y$. Hence altogether we have